Loriann, a single taxpayer, had the following income and deductions for the tax year 2018: ick he icon to view he income and deductions Click the con towe w the standard deduction amounts ick the con to view the 2018 tax rate schedule for the Single ling status Read the requirements Requirement a. Compute Lorianns taxable income and federal tax liabilty for 2018 First calculate the gross income, then calculale taxable income and the federal tax iability. (Calculate the tax using the tax rate schedule. Do not round interim tax calculations. Round the amount entered into the cell to the nearest whole dollar. If an input field is not used in the table, leave the input field empty, do not select a label or enter a zero) Total income from whatever source derived Minus: Exclusions, as provided in the tax law Gross incomeRequirements a. Compute Lorianns taxable income and federal tax liability for 2018 (round to dollars). b. Compute Lorianns marginal, average, and effective tax rates. C. For tax planning purposes, which of the three rates in Part b is the most important?Salary Business Income Interest income from taxable bonds Tax-exempt bond interest TOTAL INCOME INCOME 61,000 30,000 15,000 3,500 109,500 9,000 22,000 31,000 DEDUCTIONS: Business expenses Itemized deductions TOTAL DEDUCTIONSSTANDARD DEDUCTION Filing Status Married individuals filing joint returns and surviving spouses Heads of households Unmarried individuals (other than surviving spouses and heads of households) Married individuals filing separate returns Additional standard deduction for the aged and the blind $ 24,000 18,000 12,000 12,000 ,300 1,600 Individual who is married and surviving spouses Individual who is unmarried and not a surviving spouse Taxpayer claimed as dependent on another taxpayers returm: Greater of (1) earned income plus $350 or (2) $1,050 * These amounts are $2,600 and $3,200, respectively, for a taxpayer who is both aged and blind. Personal and Dependency Exemptions Suspended: In conjunction with the increased standard deduction amount, the Tax Cuts and Jobs Act reduces the personal exemption amount to 0 for tax years from 2018 through 2025, effectively suspending the exemptions for these years.Reference Single If taxable income is: Not over $9,525 Over $9,525 but not over $38,700 Over $38,700 but not over $82,500 Over $82,500 but not over $157,500 Over $157,500 but not over $200,000 Over $200,000 but not over $500,000 Over $500,000. The tax is: 10% of taxable income. $952.50 + 12% of the excess over $9,525. $4,453.50 22% of the excess over $38,700. $14,089.50 + 24% of the excess over $82,500. $32,089.50 + 32% of the excess over $157,500. $45,689.50 + 35% of the excess over $200,000. $150,689.50 + 37% of the excess over $500,000.
Answer #1

Part A

Salary

61000

Business income

30000

Interest income from taxable bonds

15000

Tax exempt bond interest

3500

Total income from whatever source derived

109500

Minus: exclusions provided in the tax law

Tax exempt bond interest

(3500)

Gross income

106000

Minus: deductions for adjusted gross income

Business expenses

(9000)

Adjust gross income (AGI)

97000

Minus: deductions from AGI

Itemized deductions

(22000)

Taxable income

75000

Tax liability (4453.50+(22%*(75000-38700)) =12439.50

12440

Part B

Marginal tax rate = 22%

Average tax rate= 12440/75000 = 16.59% (if answer required to be rounded to 0 decimal, it will be 17%)

Effective tax rate= 12440/(109500-9000) = 12.38% (if answer required to be rounded to 0 decimal, it will be 12%)

Part C

The marginal tax rate is the most important.

The main reason for it is that it measures tax savings possible from every additional single dollar deduction.